OpenStudy (anonymous):

i= i^2= i^3= I^4= i^5= i^6=

OpenStudy (loser66):

your opinion?

OpenStudy (anonymous):

what do they =

OpenStudy (loser66):

it is equal

OpenStudy (anonymous):

i^2=-1 i^4=1

OpenStudy (anonymous):

??

OpenStudy (loser66):

yes, you are right

OpenStudy (anonymous):

whats i

OpenStudy (anonymous):

1

OpenStudy (anonymous):

??

OpenStudy (anonymous):

i is the square root of -1.

OpenStudy (loser66):

http://www.purplemath.com/modules/complex.htm

OpenStudy (anonymous):

so what would i^5 be

OpenStudy (anonymous):

would it be i

OpenStudy (xapproachesinfinity):

\(i^5=i^4i=(1)i=i\)

OpenStudy (anonymous):

ok thanks

OpenStudy (xapproachesinfinity):

yw